OpenStudy (anonymous):

Simplify Square root of 50 + square root of 128.

OpenStudy (cruffo):

Note that 50 = 25*2, so \[\sqrt{50} = 5\sqrt{2}\] Also, 128 = 64*2, so \[\sqrt{128} = 8\sqrt{2}\] Thus, \[\sqrt{50} + \sqrt{128}\] \[= 5\sqrt{2} + 8\sqrt{2}\] Gathering like terms, \[= 13\sqrt{2} \]